How many grams of CaCO3 (Ksp = 5.0 × 10–9) is soluble in 500.0 mL of water at 25 °C.


1
Expert's answer
2021-12-07T07:44:03-0500

Solution:

CaCO3(s) ⇌ Ca2+(aq) + CO32-(aq)

The Ksp expression for CaCO3(s) is:

Ksp = [Ca2+][CO32-] = 5.0×10–9


CaCO3(s) ⇌ Ca2+(aq) + CO32-(aq)

S(CaCO3) = [Ca2+] = [CO32-] = S

Ksp = [Ca2+][CO32-] = [S] × [S] = S2


Calculate solubility (S) of CaCO3(s):

S = (Ksp)0.5 = (5.0×10–9)0.5 = 7.07×10–5

S(CaCO3) = 7.07×10–5 mol/L


The molar mass of CaCO3 is 100.087 g/mol

V = 500.0 mL = 0.5 L


Calculate the mass of CaCO3:

m(CaCO3) = S(CaCO3) × M(CaCO3) × V

m(CaCO3) = (7.07×10–5 mol/L) × (100.087 g/mol) × (0.5 L) = 0.00354 g CaCO3 = 3.54×10–3 g CaCO3

m(CaCO3) = 3.54×10–3 g


Answer: 3.54×10–3 grams of CaCO3